# In Praise of Stubbornness: The Case for Cognitive-Dissonance-Aware Knowledge Updates in LLMs

### v1

@@ 2 Dissonance-aware Targeted Updates
Figure 2: Early demonstration of a fundamental pattern we discovered across model sizes: the catastrophic nature of dissonant updates compared to non-dissonant ones. Results shown for GPT-2 Small 2(a), GPT-2 XL 2(b), and GPT-J-6B 2(c), comparing full fine-tuning (stars) and LoRA (crosses) approaches. The stark contrast between dissonant (red) and non-dissonant (green) updates persists across model scales and training methods [^3] For full finetuning, models trained until convergence on new facts, while LoRA experiments used fixed hyperparameters across both conditions. This dual approach revealed two phenomena: (1) not shown in the figure, full finetuning needed twice as many epochs to learn dissonant information compared to non-dissonant and (2) under fixed conditions with LoRA, unlike non-dissonant facts, models struggled to learn dissonant information (lower y-axis values for red crosses) while still exhibiting the same catastrophic interference with existing knowledge (low x-axis values)., motivating our systematic investigation into this fundamental challenge of knowledge integration.
2.1 Extraction of historical activations and gradients
We maintain an aggregate profile of neuronal activity by accumulating activations and gradients for each neuron at every training step. Specifically, for each neuron $n$ in the Transformer blocks—including feed-forward (MLP) layers and attention projections (Key, Query, Value matrices)—we compute $H\hat{G}_{n}$ , the cumulative historical gradient magnitude over time, and $H\hat{A}_{n}$ , the cumulative historical activation magnitude over time. To mitigate scale differences across layers, we also experiment with layer-wise normalization of activations and gradients before accumulation. Precise notation and more details are in Appendix B.
This historical activity data enables us to classify neurons as “plastic” or “stubborn” based on their past usage (see Fig.3 for a visual illustration), which is useful for our targeted network updates. We use the historical data also to normalize the input features when classifying facts as we see next.
2.2 Dissonance and Novelty Awareness
We cast our classification problem on three classes: for a given input sequence $X$, decide if it is Novel (e.g. could be integrated), Familiar (e.g. can be ignored), or Dissonant (likely requiring proper resolution).
We design a simple classifier that leverages activation and gradient information to assess the nature of new information. For any input sequence $X$, we first perform a forward pass to obtain its current activations and a backward pass to obtain its current gradients (without updating the model weights). Since the goal is to assess feasibility using easy-to-compute features and lightweight methods that could be integrated into large-scale models, we extract for each layer the mean, standard deviation, minimum, maximum, and quartiles (Q1, Q2, Q3) of the activations and gradients, eventually first normalized by historical activations and gradients. We perform ablation studies to assess the importance of different features and employ feature importance analyses to understand which aspects contribute most to the classifier’s performance. We evaluate our ability to classify facts in Sec. 3.2. Despite using simple classifiers like Random Forests and SVMs, we achieve high accuracy, opening the way for future integration of dissonance awareness into LLM training pipelines.
Refer to caption Refer to caption
Figure 3: Illustration of how we use historical neuron activity (here we show the distribution of cumulative gradients during a GPT-2-xl previous training) to identify localized areas where to store future knowledge, according to four strategies.
2.3 Targeted Neuron Updates
Building upon the historical tracking of neural activity, we implement targeted network updates to study the impact of knowledge placement location on new knowledge ingestion and past knowledge retention. We design four main types of targeted updates, which we experimentally evaluate. During training on new information, we perform standard forward and backward passes to compute the loss and gradients. Before the optimizer step, we modify the gradients to freeze certain neurons. Specifically, given the gradients for all parameters of a given layer, we zero-out those that do not belong to the selected set of neuron and corresponding weights, defined as plastic, stubborn, candidate and specific, as described below. This process effectively freezes the weights of non-selected neurons, allowing for targeted updates to specific parts of the model. By varying the choice of selected neurons, we control how new information is integrated into the model while managing its impact on existing knowledge. Next, we introduce strategies to select which neurons and weights to update: Figure 3 illustrates the conceptual relationship between the various neuron updates strategies within the model’s parameter space.
Plastic Neurons. Neurons underutilized during past model updates. To identify them, we rank neurons by increasing historical gradient values and select the top $N$ neurons with the lowest cumulative gradients:
$\mathcal{N}_{\text{plastic}}=\{n\mid\text{rank}(H\hat{G}_{n})\leq N\},$
where $H\hat{G}_{n}$ is the historical gradient for neuron $n$, accumulated over all prior training. This allows to assess whether targeting underutilized neurons can integrate new knowledge without interfereing with existing one.
Stubborn Neurons. Neurons that accumulated high historical gradients, indicating significant involvement in previous learning. We rank neurons by decreasing historical gradient values and select the top $N$ neurons:
where $|\mathcal{N}|$ is the total number of neurons, and $H\hat{G}_{n}$ is the historical gradient for neuron $n$. Updating stubborn neurons allows us to test the model’s capacity for knowledge integration and assess the potential risks of overwriting existing information.
Candidate Neurons. These neurons are relevant for encoding new information: to identify them, we perform a single back-propagation pass on the new input data, without updating the model weights. We then rank neurons based on the magnitude of these gradients and select the top $N$:
where $G_{n}^{\text{new}}$ is the gradient for neuron $n$ obtained from the back-propagation pass on the new input data. Targeting candidate neurons focuses updates on areas of the network that are most relevant to the new information, as suggested by the back-propagation process.
Specific Neurons. In another strategy, we identify neurons from the candidate set that do not intersect with the stubborn, in an attempt to store new information efficiently while avoiding interference with existing knowledge. For this, we first: (1) identify stubborn neurons $\mathcal{N}_{\text{stubborn}}$, using $N$ as defined earlier; we next (2) rank all neurons based on the magnitude of their gradients $G_{n}^{\text{new}}$ obtained from a single back-propagation pass on the new data, without updating model weights; finally, (3) we select few specific neurons, by choosing the top $N$ neurons that are not in $\mathcal{N}_{\text{stubborn}}$:
where $\mathcal{N}_{\text{all}}$ is the set of all neurons ranked by their gradient magnitudes. This last approach ensures that we select neurons that are most relevant to the new information (high gradient) while explicitly avoiding those that are crucial for existing knowledge (stubborn neurons).
3 Experimental evaluation
We discuss our experimental setup, to evaluate dissonance-awareness, as well as continual knowledge update.
3.1 Experimental setup
Dataset.  We use the COUNTERFACT dataset (Meng et al., 2022b) as our primary data source as it contains both facts and counterfacts. [^4] While this dataset allows us to test models’ handling of conflicting knowledge, we acknowledge its limitations in representing more complex real-world knowledge, a limitation which we plan to address in the future. This dataset, with approximately 17,000 facts, allows us to test models’ handling of conflicting knowledge and addition of potentially known information, [^5] For instance, general facts that pre-trained models likely were exposed to during training. two key aspects of our dissonance-aware approach. To address the lack of truly novel facts in COUNTERFACT, we generate additional data using GPT-3.5. We transform existing statements into plausible yet fictitious information, maintaining structural similarity while introducing novel content. For example, “Danielle Darrieux’s mother tongue is French” becomes “Sylvan Myrthil’s mother tongue is Sylvan” (see Appendix C.1 for details).
For our dissonance awareness experiments, we construct a balanced dataset comprising 1,000 samples each of familiar, conflicting, and novel facts. When using the pre-trained GPT-2-small model, we adjust the familiar class to 600 samples due to the limited number of known facts extracted from the model’s pre-training. For our targeted update experiments, we use 5-fold cross-validation varying each time the sets of old and new facts but keeping the following proportions: 2000 old facts vs. 1000 new facts. For conflicting updates, we also test with 10 and 100 new facts.
Models.  We employ models from the GPT family to ease comparison, focusing mostly on GPT-2-small and GPT-2-xl (and to lesser extent GPT-J 6B). The dataset size ($\simeq$17,000 facts) limits full stress-testing of larger models like GPT-2-xl (less visible catastrophic forgetting in case of non-dissonant updates compared to compressed models). As a result, some of the effects of our experiments are most clearly observed with GPT-2-small, on which we focus most in the main body of this paper, deferring GPT-2-xl results (which are aligned to GPT-2-small) to the Appendix.
We implement experiments using Hugging Face Transformers on NVIDIA GPUs. Before setting learning rates and epochs, we conduct a search for optimal hyperparameters that allow effective learning of facts (see App. D.2 for an example for GPT-2-xl). We perform the search based on the ability to correctly learn 10,000 facts from the dataset. More details including results can be found in the code repository.
3.2 Dissonance awareness
Settings.   Our first goal is to evaluate the ability to discriminate familiar, novel, and conflicting information using the readily-available [^6] We explore the use of model output-only features in Appendix. C.5 showing that using output probabilities as feature is also successful. simple features we extract from the models during the forward and backward passes. As schematized in Fig. 1 (left), we do so by relying on simple classifiers (random forests and SVMs), contrasting two scenarios for the input features: (1) a GPT-2 model fine-tuned on 1000 facts (the knowns), and (2) a GPT-2 pre-trained model (using its 600 extracted known facts as known class samples).
For each scenario, we compile a balanced dataset with equal examples per class (familiar, novel, conflicting). To create novel facts, we employ GPT-4 with carefully designed prompts, using the structure of known facts (subject, relation, object) as templates, replacing key elements with fictitious (but plausible) information. This method ensured structural similarity to known facts while maintaining novelty. Appendix C.1 provides detailed prompts and examples (full datasets will be made available upon acceptance).
@@ 3.2 Dissonance awareness
Classification performance   We extracted activations (A) and gradients (G) as described in Appendix B, and experimenting with A, G and A+G as input feature sets, using raw (R), per-layer (L) and historical (H) normalization strategies. As classifiers, we employ Random Forest (RF) and Support Vector Machines (SVM), optimizing hyperparameters using Bayesian search with 5-fold cross-validation. For clarity, we report the best results for each combination in Table 2 (average and standard deviation accuracy over the 5-folds) and defer the full results and ablation study to Table 4 in the appendix for the interested reader.
@@
Models consistently achieve high performance. Using features from the finetuned model reaches as high as (99.5%), but also using features from a pre-trained model still achieves decent performance (94.7%). Interestingly, combining activations and gradients consistently outperformed using either feature set alone, with a slight advantage of SVM over RF. Also, historical normalization helps SVM, but does not provide benefits for RF.
@@
Feature importance  While a full analysis of feature explainability is outside the scope of this paper, we further seek to observe structural differences in the learning process by comparing feature importance, using the feature importance scores derived from the random forest algorithm.
To further shed light on such difference, Fig. 4 opposes feature importance in both cases, focusing on Activation versus Gradient-related features. It turns out that in the finetuned scenario, gradient-based features are substantially more important. This is likely due to the fact that finetuning the models on these facts has somewhat overfit them leading to gradients that are more discriminative: e.g. a clearly null gradient for known facts and a clearly high one for unknown ones. For the pretrained scenario, however, which is the most likely case in a real case scenario, both activation and gradient features contribute significantly, suggesting that for long-term knowledge, both internal representations and learning dynamics should be mixed in order to achieve good classification. Appendix.C.3 expands this analysis by focusing on transformer block importance instead.
@@
Key findings   Overall, despite the simplicity of our features, the results demonstrate the feasibility of distinguishing between familiar, novel, and conflicting information, even in the challenging case of using pre-trained models, providing the needed foundation for dissonance-aware updates, which we explore in the next experiments.
3.3 Non-dissonant updates
Settings.   We now investigate how LLMs handle non-dissonant updates using our different strategies as experimental tools. In our experiments, schematized in Fig. 1 (middle), we evaluate the incorporation of non-conflicting facts into our models. The pipeline consists of (i) training on 2,000 initial facts (old) while collecting historical gradients to identify stubborn neurons, (ii) simulating updates with 1,000 new facts (new) to collect current gradients for candidate identification and (iii) applying different neuron selection strategies to update the model with these 1,000 facts. Note that while we track 2,000 facts as proxy for old knowledge, this represents a smaller fraction of GPT-2-xl’s total knowledge compared to GPT-2-small, limiting our visibility into effects on other untracked pre-trained knowledge.
For our experiment below inspired by the lottery ticket hypothesis (Frankle & Carbin, 2018), we used 10,000 separate facts for gradient extraction before training a fresh model on the 2,000+1,000 facts described above.
Results.   Fig. 5 presents the accuracy of various neuron update strategies on old and new knowledge for GPT-2-small, including error bars representing standard deviations over five runs. We observe that simple fine-tuning leads to a slight degradation in performance on old knowledge, dropping to approximately 93% accuracy. In contrast, updating plastic neurons helps preserving old knowledge, with accuracy remaining above 98% even when using up to 20,000 neurons. Random neuron selection exhibits a similar behavior. However, using candidate or stubborn neurons results in slightly more degradation to old knowledge.
(a) Old Knowledge
(b) New Knowledge
Figure 5: Non-dissonant updates: Old vs new knowledge for targeted updates on GPT-2-small (see Fig. 10 and 11 for GPT-2-XL)
Interestingly, the strategies differ also in their ability to ingest new knowledge. Plastic neurons preserve old knowledge but struggle to ingest new knowledge requiring comparatively more neurons. This contrasts with specific neurons which strike the best balance, learning new knowledge with fewer neurons while maintaining acceptable levels of old knowledge retention.
An intriguing pattern is that targeting stubborn, candidate, or specific neurons allows the model to learn new knowledge using fewer parameters compared to targeting plastic neurons. This finding resonates with the existence of winning subnetworks, as suggested by the lottery ticket hypothesis (Frankle & Carbin, 2018). It implies that certain subnetworks within the model are more conducive to integrating new information, compared to others. We conduct further experiments that confirm this hypothesis in App D.1.
Finally, we conducted the same experiments on GPT-2-xl, where our 2,000 tracked facts represent a much smaller portion of the model’s knowledge. With this larger capacity, interference with tracked facts becomes naturally less likely, leading to all strategies showing even better preservation of our monitored knowledge, compared to GPT-2-small. As shown in Fig. 11 and Fig. 10 in the appendix, all strategies tend to preserve old unrelated knowledge; and differ, as in the case of GPT-2-small, in their ability to integrate new knowledge efficiently. Detailed analyses, including the effects of varying learning rates and neuron counts, are provided in Appendix D.3.
Key Findings.  LLMs show remarkable robustness when incorporating non-dissonant information, as long as heavily-used (stubborn) neurons are avoided. Updating plastic neurons helps preserving old knowledge but requires more parameters (or time) to achieve high accuracy on new knowledge. Targeting specific neurons offers a balanced approach, enabling efficient knowledge integration with minimal impact on existing information.
3.4 Dissonant Updates
Settings.   We examine how LLMs handle conflicting information as shown in Fig. 1 (right): after training on 2,000 old facts and 1,000 new facts, we introduce 1,000 conflicting updates contradicting previously learned new facts, applying our various strategies to study their impact on learning conflicting facts and preserving unrelated old knowledge.
Results.   We illustrate the performance of various strategies on GPT-2-small in Fig. 6. Again, old knowledge refers to the 2000 initial completely unrelated facts and new knowledge corresponds to the 1000 conflicting ones. Surprisingly, we find that dissonant updates are highly destructive to the retention of such unrelated knowledge, regardless of the neuron update strategy employed. Even when updating plastic neurons, which are presumed to be underutilized and thus less likely to interfere with existing knowledge, we observe significant degradation in the model’s ability to recall old facts.
(a) Old Knowledge
(b) New Knowledge
Figure 6: Dissonant updates: Impact of 1000 dissonant facts and GPT-2-small (see Fig. 13 for GPT-2-XL)
Given the destructive nature of simultaneously adding 1,000 conflicting facts, we conducted additional experiments where we added only 100 and 10 facts. The results, detailed in Appendix E.1, indicate that while the impact on old knowledge retention is less severe when editing fewer facts, the destructive effect remains prominent. This reminds us the performance of state-of-the-art model editing methods such as ROME (Meng et al., 2022a) and MEMIT (Meng et al., 2022c) which also deteriorates when applied to multiple sequential edits, as opposed to single-edit evaluations typically reported in the literature. While our primary focus is not on developing new model editing techniques, we leverage EasyEdit (Wang et al., 2023) to benchmark the above existing methods under our same multi-fact experimental conditions (See App. E.2).
Finally, we also performed experiments with GPT-2-xl under various conditions, deferred to Appendix E.4 for space constraints. Overall, similarly to the non-dissonant case, GPT-2-xl exhibits different scaling behaviors. Surprisingly though, even when not learning conflicting knowledge, and despite having much more parameters, GPT-2-xl also experiences significant degradation in old knowledge retention– further confirming the catastrophic nature of dissonant updates (See Fig. 13). As shown in Fig. 3, GPT-J6B also confirms the same tendency.
Key Findings.  Dissonant updates pose a significant challenge, as they are destructive to prior unrelated knowledge, regardless of model size and even when targeting unused neurons. This underscores the importance of dissonance awareness to detect and appropriately handle conflicting information during continual learning. Our results motivate the integration of dissonance classifiers directly into the update or training of large language models. Thus, developing dedicated conflict resolution methods remains an essential direction for future work.
@@ 4 Discussion and conclusions
4.1 Lessons learned
Fundamental Properties of Knowledge Updates: Our results reveal a striking pattern: while non-dissonant updates show remarkable robustness, dissonant updates trigger severe corruption of unrelated knowledge - dropping accuracy below 60% even when modifying just 10-100 facts. This effect persists across our tested model scales and update strategies, suggesting a deeper challenge in how current neural architectures handle contradictory information.
Feasibility of Dissonance Detection: LLMs encode clear signatures distinguishing between novel, familiar, and dissonant information. In our controlled experiments, simple classifiers achieve 95% accuracy with pre-trained models and 99% with finetuned models using either activation/gradient features or output probabilities. This suggests the potential for developing mechanisms to identify potentially problematic updates before they occur.
Promise of differentiated plasticity: Avoiding heavily-used neurons during non-dissonant updates improves robustness, maintaining 98% accuracy on old knowledge (versus 93% with finetuning). Interestingly, neurons used during pre-training are particularly effective at integrating new knowledge, extending lottery ticket hypothesis findings (Frankle & Carbin, 2018) to language models.
4.2 Limitations and Future Directions
Experimental Control vs. Scale: While our controlled experiments reveal fundamental properties of knowledge updating, investigating these phenomena in much larger models presents challenges as it is not straightforward to track the impact on their broader knowledge.
Dataset Limitations: Our current findings rely on CounterFact-derived data with simple factual statements. Developing larger, more diverse datasets is essential for understanding how these properties generalize to more complex forms of knowledge and conflicts.
Neuron Classification Metrics: While our analysis of neural plasticity uses gradient magnitudes effectively, future work could explore richer metrics incorporating activation patterns and network connectivity to better understand knowledge distribution and update mechanisms.
Beyond Binary Dissonance: Our current investigation treats dissonance as binary, while real-world knowledge updates often involve varying degrees of conflict and different types of knowledge. Understanding how these nuances affect knowledge integration remains a challenge.
Towards Human-Inspired Updates: The catastrophic nature of dissonant updates suggests we may need fundamentally different approaches to LLM training. Rather than attempting to overwrite existing knowledge, future work might explore mechanisms for maintaining and contextualizing potentially conflicting information - similar to how humans maintain both historical and updated knowledge with appropriate contexts.
Impact Statement
These findings motivate our development of dissonance detection capabilities as a crucial safety mechanism for deployed AI systems. More broadly, they suggest we may need to fundamentally rethink AI architectures to develop systems that, like humans, maintain and contextualize potentially conflicting information rather than attempting to overwrite it. This might require moving away from current approaches that try to ”edit” neural networks toward architectures that can accumulate and contextualize knowledge while maintaining multiple, temporally-organized versions of truth.

### v2

@@ 3 A tale of two updates: one safe, the other catastrophic
Figure 3: Safe Non-dissonant vs. Catastrophic Dissonant Updates. Results shown in one of our folds, for GPT-2 Small 3(a), GPT-2 XL 3(b), and GPT-J-6B 3(c), comparing full fine-tuning (stars) and LoRA (crosses) approaches. The stark contrast between dissonant (red) and non-dissonant (green) updates persists across facts, model scales and training methods.
@@
Note that for full finetuning, models were trained, as per our protocol, until convergence on new facts, while LoRA experiments used fixed hyperparameters across both dissonant and non-dissonant conditions. This dual approach allows us to illustrate two phenomena: (1) not shown in the figure, full finetuning needed twice as many epochs to learn dissonant information compared to non-dissonant one and (2) under fixed conditions with LoRA, unlike non-dissonant facts, models struggled to learn dissonant information (lower y-axis values for red crosses) while still exhibiting the same catastrophic interference with existing knowledge (low x-axis values). This suggests that contradictions force dramatic reorganization of the model’s weight space, disrupting unrelated knowledge, while non-dissonant facts integrate naturally into existing structures.
4 Selective Plasticity: Helpful for One, Hopeless for the Other
Our investigation of selective plasticity reveals another fundamental asymmetry: while avoiding heavily-used neurons successfully protects knowledge during non-dissonant updates, no targeting strategy can prevent catastrophic interference during dissonant updates.
The Asymmetry of selective plasticity Fig. 4 presents the stark contrast between non-dissonant (top row) and dissonant (bottom row) updates using various neuron targeting strategies on GPT-2-small (with error bars representing standard deviations over five runs). While all strategies show relatively safe behavior for non-dissonant updates, they all prove catastrophic for dissonant ones.
Non-dissonant Updates: Selective Plasticity Works. When incorporating new, non-contradictory information, avoiding stubborn neurons dramatically improves knowledge preservation. Standard fine-tuning drops old knowledge retention to  93%, but targeting plastic neurons maintains >98% accuracy even when updating 20,000 neurons. Random selection achieves similar protection, likely by avoiding stubborn neurons “by chance". However, targeting candidate or stubborn neurons leads to more degradation of existing knowledge.
Dissonant Updates: All Strategies Fail. The picture reverses completely for contradictory information. Even targeting plastic neurons can result in worse retention than standard fine-tuning. Dissonant updates prove catastrophically destructive regardless of neuron selection strategy, destroying unrelated knowledge across all approaches.
(a) Non-dissonant: Old Unrelated Knowledge
(b) Non-dissonant: New Knowledge
(c) Dissonant: Old Unrelated Knowledge
(d) Dissonant: New Knowledge
Figure 4: The Selective Plasticity Asymmetry (GPT-2-small). While avoiding stubborn neurons preserves old knowledge during non-dissonant updates, all strategies fail catastrophically with dissonant updates. See Fig. 9 and 12 for GPT-2-XL results.
Existence of preferred Subnetworks Interestingly, across both update types, we observe that targeting stubborn, candidate, or specific neurons enables more efficient learning of new knowledge compared to plastic or random neurons. This finding resonates with the existence of winning subnetworks, as suggested by the lottery ticket hypothesis (Frankle and Carbin, 2018). It implies that certain subnetworks within the model are more conducive to integrating new information, compared to others. We conduct further experiments that confirm this hypothesis in App D.1.
Scale Invariance of the Asymmetry. Experiments with GPT-2-XL confirm the same asymmetric pattern, though with different visibility and learning dynamics due to scale. As mentioned in our experimental pipeline, our 2,000 tracked facts represent a smaller fraction of GPT-2-XL’s knowledge, making interference less observable (though not absent). For non-dissonant updates, all strategies preserve monitored knowledge even better than in GPT-2-small, while maintaining the key finding that avoiding stubborn neurons provides better protection (see Fig. 10 and Fig. 9). Critically, dissonant updates remain catastrophic even when targetting plastic neurons, confirming this as a fundamental limitation rather than a capacity issue (see Fig. 12).
Robustness Across Contradiction Scales. Finally, to test whether the catastrophic effect scales with the number of contradictions, we conducted experiments with 10, and 100 dissonant facts only. While smaller contradiction sets show slightly less severe effects for targeted strategies, the destructive impact remains prominent across all scales. Standard fine-tuning with just 10-100 contradictions can still corrupt up to 80% of unrelated knowledge (detailed results in Appendix E.1).
5 Detecting Dissonance Before It Causes Damage
We now investigate the feasibility of dissonance detection using our classification task as a proxy. For lack of space, we report results for the internal features and defer the reader to Appendix C.5 for output model features, which achieved equally good performance.
@@ 5 Detecting Dissonance Before It Causes Damage
Classification performance   We use combinations of activations (A) and gradients (G) as described in Appendix B as input features, using raw (R), per-layer (L) and historical (H) normalization strategies. We report the best results for each combination in Table 2 (average and standard deviation accuracy over the 5-folds) and defer the full results and ablation study to Table 4.
@@
Using features from the finetuned model, we reach as high as 99.5%, but also using pre-trained model features still achieves decent performance (94.7%). Not shown, combining activations and gradients consistently outperformed using either feature set alone, with a slight advantage of SVM over RF.
@@
Feature importance  While a full analysis of feature explainability is outside the scope of this paper, we compare feature importance between the two scenarios, using the scores derived from the random forest algorithm. Fig. 5 reports the results, focusing on Activation versus Gradient-related features. It turns out that in the finetuned scenario, gradient-based features are substantially more important. This is likely due to the fact that finetuning the models on these facts has somewhat overfit them leading to gradients that are more discriminative: e.g. a clearly null gradient for known facts and a clearly high one for unknown ones. For the pretrained scenario, however, which is the most likely case in a real case scenario, both activation and gradient features contribute significantly. Appendix.C.3 expands this analysis by focusing on transformer block importance instead.
@@
Takeaway  Using either internal features (as we show here) or output features (appendix), dissonance detection is feasible on our COUNTERFACT-augmented classification dataset, offering hope for early detection and prevention of catastrophic interference.
@@ 6 Discussion and conclusions
Through systematic empirical investigation, we uncovered a fundamental and previously unknown asymmetry in how language models handle updates: while non-contradictory updates are integrated safely, contradictory updates trigger catastrophic corruption, destroying up to 80% of unrelated knowledge, even with as few as 10-100 contradictory facts. This effect persists across model scales and training approaches, suggesting a fundamental limitation rather than a capacity issue.
Our findings seriously question the current practice of training, continual learning, and knowledge editing in LLMs. The very existence of updates that catastrophically impact unrelated knowledge not only reveals a critical vulnerability in current architectures but could potentially serve as an attack vector in deployed systems. This vulnerability becomes particularly concerning in real-world applications where systems inevitably encounter contradictory information, whether through evolving scientific understanding or potential (malicious) misinformation campaigns.
Our work further offers inspiration for understanding continual learning in biological systems by highlighting a crucial pattern we all know but hadn’t formally recognized: the power of append-only updates that never overwrite, as when humans naturally maintain that “Pluto was once classified as a planet; today it is not.” Our empirical findings with artificial neural networks offer further inspiration, demonstrating that selective plasticity indeed helps with non-contradictory updates while failing with contradictory ones, suggesting that avoiding overwrites, through episodic (here time) contextualization, might be fundamental to robust knowledge accumulation.
Finally, our demonstration that contradictions can be detected with high accuracy (95%+) using simple model features offers hope for protective mechanisms. While this initial investigation requires scaling to more complex datasets and larger models (See limiations in Sec. F), it suggests that early detection and prevention of catastrophic interference may be feasible. These findings open crucial new research directions in continual learning, model editing, and the development of more robust AI systems that can, like biological systems, safely accumulate and contextualize potential conflicts.